A machine has a 16-bit byte-addressable virtual address space. The page size is 8 KB. How...

  1. A machine has a 16-bit byte-addressable virtual address space. The page size is 8 KB. How many pages of virtual address space exist? (25 points)
Answer #1

Given:
Virtual Address size of the machine: 16bits
Page Size : 8Kb

Thus, before calculating each address is to converted into the number of bytes

Therefore
Virtual address size = 2^16 bytes
Page Size = 2^13 bytes

Now using the formula

Number of entries in page table = (virtual address space size)/(page size)

Number of pages = (2^16)/(2^13) = 2^3

Thus,
Number of pages of virtual address space = 2^3